Came into lab today, with two main goals:
1) Bring Downs equipment for vibration measurements of cryo cooler during next operation
--> DONE, see photo
2) Assemble shields, with two layer scheme using electropolished and mill finish panels.
--> issue: the electropolished shield set was only partially shipped out in Ticket 15101 - unfortunately there was a misinterpretation by Logistics personnel regarding which parts to ship, due to the formatting of the ticket page. I will coordinate shipment of the balance of the panels to the vendor Able Electropolishing for completion of PO S479514
I entered QIL just before Wed Sep 23 00:27:51 2020 to check out and photograph the sprinklers, spent about 20 min looking around the lab and drawing inspiration for in cryo. Wore shoe covers and gloves, touched nothing, sanitized doors after.
HEPA filters on top of the WOPO table have been turned to High (earlier were at Low).
2020 Sep 01, StephenA with remote assistance from RaymondR
Cryopump is up and running. Initial attempts to run the cooler were stymied by an open circuit in the cold head to compressor connection caused by one of the two accessory port fuses (right, circled in attachment 1). The compressor would run but the valve motor wouldn't start in the cooler itself. I extended the spring in the fuse housing (attachment 2) and it seems to have fixed the problem, as now the valve motor starts at the same time the compressor is turned on. Attachment 1 also shows the highly technical cord management procedure done to reduce the trip hazard caused by the compressor power plug.
User manual recommendations*:
*Manuals for both the compressor and the cryocooler are linked on the West Bridge wiki manuals page
The diaphragm pump was turned on earlier this week after finally closing up this external adapter tank. Out of an overabundance of caution the tank and cryocooler are supported by the skycrane and a number of posts to prevent it walking off the foam resting pad once the cryocooler is switched on.
All temperature sensors agree with each other within 0.1 K at room temperature
Used the 19-pin MIL feedthrough to run 4 platinum RTD's and a 25 Ω 100 W resistive heater to the cold head. Attachment 1 is the wiring diagram for the feedthrough and the D-sub connector to the CTC-100 temperature sensor. Attachment 2 shows the three RTDs placed on the cold head. It also shows the thermal anchoring of all lead wires. Attachment 3 shows the RTD attached to the cooler below the cold head using cigarette paper and cryo varnish (stored in the flammables cabinet in QIL).
The Al block is a premade PT-RTD integrated mounting setup, which was placed on some indium sheet bits and clamped down with a screw and belleville washer. The other two cold head sensors are pressure fit to the cold head by a spring loaded mini dog clamps, and one of the two has some indium underneath the RTD to see if there is any value in doing so going forward with these mounting springs. The glued sensor was attached by painting a thin layer of cryo varnish on the cooler, adding a strip of cig paper, layer of varnish, press in sensor, another strip of paper, paint over all of it with a last thin layer of varnish that reaches beyond the bounds of the paper strips.
Picked up the prototype shield panels from Hamilton Metalcraft 7/22 and brought them to QIL. All of the parts are wrapped by part number and in a bin (see attached photo). There are 6 sets of shield panels, but 2 full sets were removed for coating vendors. One full set is as follows (20 parts total):
All component #'s are preceded by 'D2000298-'. 031, 032, and 033 are 03 panels but with hole variations, same goes for 131, 132, and 133 with respect to panel 13
Alex dropped off the new round of 2um PD's, they're on the north table accompanied by his data sheet.
First day back (7/15) found the particle board trim w/ powerstrip on the QIL workbench had collapsed. Re-glued and added 4 screws to the middle board where vertical boards from the shelves extended low enough. See attached photos for before, during, and after looks.
Copper parts picked up July 23rd and brought to QIL, now only waiting on PO# S477874 and the pirani gauge from Koji's bulk JPL order
Torque driver set for QIL setup bolted joints, with range 15 in*oz - 50 in*lb, p/n WIHA 5HYL9, is on order from Grainger, with anticipated delivery in the week of July 20th. Refer to PO S477925. *update* Tracking Number UPS 1Z19W9330321365493
Cryo connection copper parts PO S475316 will be finished early next week by the machine shop in Torrance, I'll bring them to campus or to Raymond's place (TBD).
I've updated the PSOMA optical layout. I still have some questions on locking, and there are a few additional configurations that we could try. In particular:
Each of these configurations also has a couple different ways to pickoff an LO for homodyne readout. Shruti and I enumerated these configurations on a zoom whiteboard a couple weeks ago, and I've attached them (the zip contains png).
Chris also mentioned last week that we may run into a frequency-dependent loss in the critically coupled cavity configurations. The pdf I've attached shows a configuration that I think is a minimal modification of the Mach-Zehnder amplifier described in PSOMA. One of the ring cavities is replaced with a tunable steering mirror, and the LO is picked off before the pump reaches the MZ.
In the new diagram, I'm thinking about controlling the following degrees of freedom:
Some things I'm unsure about:
Shruti and I are now tracking our work on git issues in the PSOMA repo.
Attachment 1 contains the SR785 dark noise measurements at number of PD reverse bias voltages from 77-295K with filenaming convention:
[PD]_drkspec_[date]_[temp]_[input V]_[scan freq]_[FEMTO gain]_[date]_.txt
It also contains the keithley sweeps for QE calculations.
I'm still working out what is wrong with the QE data and how to effectively process the dark noise versus temperature.
I put together two PSOMA layouts, one for a bowtie cavity and one for a ring cavity configuration.
I expect there are a number of problems with the layout as I've drawn it, and I note a number of these in the bowtie diagram. Among these
I did not see anyone in the building.
Attachment 1/2: Our labs have no sticker/paper to indicate any disinfection of the room. (Make sense)
Attachment 3: Most of the basement offices have the notes to indicate disinfection.
Attachment 4/5: Our offices have no notes.
To check the status of all the labs, I went to WB. There was no ongoing water leakage in the labs.
Attachment 1: The subbasement was completely dry.
Attachment 2: Upon the lab inspection, I took PPE from the OMC lab. This was intended to prevent me to pick up anyone's anything and you to pick up my anything.
Attachment 3: The EE shop has no problem
Attachment 4: Cryo Lab. No problem.
Attachment 5: Crackle Lab. No problem, but a lot of dead cockroaches on the floor!
Attachment 6: OMC Lab. No problem.
Attachment 7: C.Ri.Me Lab. Gabriele has already checked the status in the morning. And I found no problem. Didn't bother to turn on the light.
Attachment 8: CTN Lab. No problem.
Attachment 9: QIL Lab. The floor was mostly dry. Did someone wipe the floor?
Attachment 10: Some water drip was found in front of the workbench.
Attachment 11: It comes from the ceiling.
Attachment 12: Left a trash box to catch future possible leak.
Attachment 13/14: TCS Lab. No problem found.
Attachment 15: As per Aidan's request, the instruments were moved to the North-East area of the room to avoid future possible leak.
West Bridge flooding Apr 6th due to rain in the night
Looks like the first responder was Calum. The attached photos were sent from him.
Last Friday, I found the HEPA units on the squeezer table were not on. I turned them on at "SLOW".
Still trying to figure out how to set up the particle counter remotely. The current particle count is 576.
Note: the particle count is the number of particles detected over 0.3um size.
Item lending as per Ian's request: Particle Counter from OMC Lab to QIL
The current particle class of the room was measured to be 800.
The particle counter went back to the OMC lab on Aug 10, 2020.
Facilities workers replumbed the water lines feeding the air handler units in the QIL. Tomorrow they plan to come back for about an hour to insulate the lines. We'll keep the tables wrapped up until that's complete.
[Raymond, Chris, Koji, Chub, Aidan, Duo]
Both tables were surrounded with plastic shielding in preparation for the HVAC maintenance/repairs to be performed tomorrow, Tuesday 24/02/2020. The IR Lab cryostat roughing and turbo pumps were turned off to avoid overheating while inside their new plastic cocoon. Sticky floor mats were placed in the oil-slick area in front of the flammables cabinet to improve safety for all foot traffic in the area, as well as to mitigate the urge to boogie while working in the lab.
Here is the data for the last week. The temperature for the Northside is noticeably higher than the Southside. This is probably a calibration error in the North sensor because there is not a noticeable temperature difference when walking across the room. I would guess the south sensor is more accurate in its overall temperature reading because it feels more like 22 degrees than 24.
Got data from the HOBO temperature recorders. The temperature has been fluctuating between about 80 degrees F to 60 degrees F. The Temperature has been fluctuating wildly in the last month that the recorder was on. I will restart the HOBO monitors to get more current data.
*Note the HOBO app changed how it reports the time data. The updated code is attached which uses the AM and PM 12 hour .csv file generated by the app.
The QEs were measured at 293K, 239K, 232K, and 293K again. The cooling was provided by the PD TEC. At each temperature, the incident power was changed from 30uW to 1mW to see the dependence of the QE on the incident power to check the possible saturation.
The QE was 79~81% (the window T=96.6% was already compensated). I'm not 100% sure this 1% variation in the plateau is real or due to insufficient calibration of the REF PD.
The REF PD was calibrated at 1mW at 100mA injection current to the laser.
No obvious saturation was observed.
We can cool the PD with LN2 and we should make a careful alignment of the beam at each temperature.
1) I've brought another TEC driver fro the PD temp control. This unit was borrowed from the 2um ECDL setup. Eventually, we need to return this to ECDL. (Attachment 1)
The PID loop of the TEC control works. But it is not well optimized yet. If you change the target temp too quickly, the TEC out seemed oscillating. Watch the TEC out carefully and change the temp setpoint slowly.
So far I have tried to cool the thermister up to 30kOhm (~232K) and I_TEC was 0.33A. I did not try further. I felt it was better to cool the PD base for further trial.
2) A part of the alignment study, the beam is aligned to A2P6. Also, the lens position was investigated, and I decided to move the lens ~1 inch away from the window. (Attachment 2)
In fact, this allowed us to insert the power meter between the lens and the window.
The QE measurements from the first couple of photodiodes are attached below.
QE = [I_photocurrent]/[P_PD] * h *nu/e
P_PD = Power incident on photodetector = 0.966*power_incident on cryo window
Power incident on cryo window = F(voltage on reference PD)
% load JPL data
f0 = dir('*dark*.txt');
f1 = dir('*photo*.txt');
f2 = dir('*cond*.txt');
% get temperature vs time
tempList = ;
pList = ;
for ii = 1:numel(f2)-1
% load JPL data
f0 = dir('*dark');
f1 = dir('*bright*');
% get temperature vs time
tempList = ;
refPDList = ;
for ii = 1:numel(f1)
Opened the cryostat to resolder the heater and re-wrap the thermal anchor for the sample RTD and PD connection. All connections are working as expected at room temperature. A2 is still in the sample mount.
Still to do:
screen -RAad autorun
don't we also want to record the dark noise spectrum as a function of T and V_Bias ? I would guess that the dark noise doesn't always scale with dark current at low frequencies since its probably more like a random walk than shot noise.
[Aidan, Chris, Ray]
[Raymond, Aidan, Chris, Koji]
P6 element (500um)^2
- We looked at the current amp (FEMTO) output. The amplifier saturated at the gain of 10^3 V/A. Looking at the output with a scope, we found that there is a huge 1.2MHz oscillation. Initially, we thought it is the amplifier oscillation. However, this oscillation is independent of the amplifier bandwidth when we tried the our-own made transimpedance amp.
- Shorting the cryostat chamber to the optical table made the 1.2MHz significantly reduced. Also, connecting the shield of the TEC/Laser controller made the oscillation almost invisible. This improvement allowed us to increase the amp-gain up to 10^7.
- Then the dominant RMS was 60Hz line. This was reduced by more grounding of the cable shields. The output was still dominated by the 60Hz line, but the gain could be increased to 10^8. This was sufficient for us to proceed to the careful measurements.
- The dark current was measured by the source meter, while the photocurrent (together with the dark current) was measured under the illumination of the ~1mW light on the PD.
- Attachment 1 shows the dependence of the dark current against the swept bias voltage. We had ~mA dark current at the room temp. So, this is ~10^5 improvement.
- Attachment 2 shows the dependence of the apparent QE against the swept bias voltage. The dark current was subtracted from the total current, to estimate the contribution of the photocurrent in the measurement.
- Attachment 3 shows the dark noise measurement at the reverse bias of ~0.6V. Up to 1kHz, the noise level was below the equivalent shotnoise level of 1mA photocurrent.
All the data and python notebook in the attached zip file.
We placed a power meter after the fiber collimator, 75mm focal length lens and HR mirror at 45 degrees - basically, we placed the power meter immediately before the input window to the cryo chamber after all the intervening optics from the fiber output.
For a series of laser diode current levels, we measured the power on the power meter and the corresponding voltage on the reference photodetector that is monitoring a 10% pick-off from the laser. The calibration is as follows:
Wow. This is great, thanks Chris.
Note: in the preceding table, channel numbers use the digital convention (numbered starting from zero), which is not the convention used by the AA/AI chassis front panel (numbered starting from one).
I can see some screws are not vented. You also need to use a vented screw for the additional temp sensor if the face screws of the PD mounts are not vented.
You can use a bunch of clean clamps and screws I brought. They are in a mylar bag.
If you need more vented screws, please specify the size and length. I can grab some from the 40m cleanroom.
The IR Labs cryostat has its internals wired and attached to the baseplate. PD A2 was clamped and the vacuum pumps turned on for the first cooling test.
[in the morning I will update with a detailed pin-out and label the attached photo (labeled 12/13, pin out in separate post)]
Recently Duo wanted to make an arbitrary waveform excitation using the QIL cymac, but it wasn't working. An excitation would die after 10 seconds or so, with awgtpman reporting that the data was too far in the future.
It turns out this was caused by a missing leap second in the RTS software. It is now fixed upstream, and we're running a patched version of awgtpman on fb4, until the change propagates to the packaged version.
I borrowed KEITHLEY 2450 source meter from Rich. The unit comes with special coaxial cables and banana clips. Most of the peripherals are evacuated in the OMC lab.
The dark current of A2P2, A2P3, A2P6 were measure with different temperatures (300K, 270K, 254K). The plot combined with the previous measurement ELOG QIL 2425.
== How to use the source meter ==
- Two-wire mode: Connect the wires to the diode
- Over voltage protection: [MENU] button -> SOURCE / SETTINGS->Over Voltage Protectiuon 2V
- Sweep setting: [MENU] button -> SOURCE / SWEEP -> e.g. Start -750mV, Stop +500mV, Step 10mV, Source Limit 1mA -> Select Generate
- Graph View: [MENU] button -> VIEWS / GRAPH
- Start measurement: Note: The response of [TRIGGER] button is not good. You need to push hard
This starts the sweep, or a menu shows up if your push is too long -> Select "Initiate ..."
- Data Saving: [MENU] button -> MEASURE / READING BUFFERS -> Save to a USB stick
Update on the 32 pin female connector:
Mouser's overnight delivery was rejected on 12/3 by mail services for being soaked in an "unkown liquid" and was therefore taken away by the courier for return to Mouser. This was the last one in stock, so I ordered a replacement through Digikey for express delivery this morning 12/4, but it has not yet arrived. I've called Fedex and the package was sorted at the LA facility but not given to the courier for morning delivery. It is now estimated for delivery this afternoon.
The quantities we want to measure as a function of the temperature:
- Temperature: 2.2k thermister resistance / 100ohm platinum RTD
- QE (Illuminating output / Dark output / Reference voltage / Reference dark output)
- Dark current (vs V_bias) -> Manual measurement or use a source meter
- Dark noise (PSD) 100kHz, 12.8k, 1.6kHz, 100Hz
[Raymond and Koji]
We dunked the PD socket test piece into LN2 and repeated heat cycle 8 times. No obvious change was observed. Then the wires were pulled to find any broken joint or etc.
None of the solder joints showed the sign of failure.
For cleanliness, we are going to use In-Ag solder (no flux) for the actual wiring.
Attachment 1: Frozen connector
Attachment 2-4: Inspection after thawing.
While I'm still waiting for the proper connector for the vacuum feedthru of the IRLabs cryostat, I have connected to the Dsub9/15 split cable to another Dsub9 connector so that I can test the cooling of the InAsSb sensor in air. Also, the 2004nm laser, a fiber-coupled faraday isolator, and 90:10 beam splitter was moved to the cryostat table and fixed on a black al breadboard. [Attachment 1]
The InAsSb TEC was controlled by the TEC controller of ITC-50. I didn't change the PID parameters of the controller but the temperature nicely setteled to the setpoint. The sensor has a 2.2kOhm thermister. And the max current for the TEC was unknown. The TEC driver had the current limiter of 0.3A and it was not changed for now. With this current limit, the thermistor resistance of 10Kohm was realized. This corresponds to the temperature of about -20degC. According to the data sheet given by Alex, the resistance/temperature conversion is given by the formula
1/T = 7.755e-4 + 3.425e-4*log(R)+1.611e-13*log(R)^3
To satisfy the curiosity, the dark current of a (500um)^2 element was measured between -250K and -300K. At -254K, the dark current went down to the level of 40uA (1/15 of the one at the room temp). For the measurement, the bias voltage was set to be 0.5 and 0.6V. However, it was dependent on the diode current. (Probably the bias circuit has the output impedance). This should be replaced by something else.
The external Dsub cable is ready except for the 32pin connector to be plugged-in to the chamber. See QIL ELOG 2460 for the pin assignment.
Normal solder (Sn63 Pb37): with flux, wetting o
Pure Indium - In 99.995: no flux, wetting x, low melting temp, like paste
Pb93.5 Sn5 Ag1.5: with flux, wetting o, high melting temp (soldering iron setting 380~430F)
Cryo solder In97 Ag3: no flux, wetting x, low melting temp, like paste
The IR Labs cryostat is now pumping down. The thor labs posts are necessary for centering the optical port at a height of 5.5". The orange cabinet used to elevate the pumping station was relocated from next to the computer (was, and still is, empty). The power supply is for the wide range vacuum gauge attached to the pumping line.
I've ordered a liquid nitrogen dewar for arrival Monday morning.
For future reference, the gauge communicates via a VGA port, for which the pin-to-wire association is as such: 1-Black; 2-Brown; 3-Red; 4-Orange; 5-Yellow; 6-Green; 7-Blue; 8-Purple; 9-Grey; 10-White; 11-Pink; 12-Cyan; 13-Black/White; 14-Brown/White; 15-Red/White (MKS 901P non-ethercat wiring diagram)
edit by RXA: replaced multi GB PDF w a reasonable JPG.
System diagram of the PD QE test with the IRLabs cryostat.
PT-SE (MS/PT-SE) connector data sheets
Amphenol catalog http://www.amphenol-industrial.com/images/catalogs/PT.pdf
Detoronics Hermeic Sealed Connectors (DT02H-18-*PN) http://www.hselectronics.com/pdf/Detoronics-Hermetic-Connectors.pdf
AF8 crimping tool (expensive!) https://www.mouser.com/ProductDetail/DMC-Tools/AF8?qs=gvhpkjpQEVSjrLbsepewjg%3D%3D
AF8 alternative https://www.jrdtools.com/?gclid=Cj0KCQiA2vjuBRCqARIsAJL5a-IQ9ztCEYKdo645v_RhUBJS3eMIars1LubjlKZoorS-lnx6ClDDiMUaAlZiEALw_wcB
Thermistor link: https://www.tec-microsystems.com//Download/Docs/Thermistors/TB04-222%205%25%20Thermistor_Specification_upd2018.pdf
TEC spec: Mounted TEC type: 2MD04-022-08/1 https://www.tec-microsystems.com/products/thermoelectric-coolers/2md04-series-thermoelectric-coolers.html
2MD04-022-08/1 dTmax = 96, Qmax = 0.4W, Imax = 0.7A, Umax = 2.0, ACR = 2.29 Ohm
The PD mounts were delivered from ProtoLabs. The order was sent on Tue last week and it's here on Monday. Excellent!
And the quality looks pretty good.
The surfaces are sandblasted. Do we want to do any process on the bottom surface to reduce the thermal resistance?
An indium solder string also came in.